OpenStudy (anonymous):

Write the first ten terms of a sequence whose first term is -10 and whose common difference is -2.

OpenStudy (midhun.madhu1987):

a = -10 d = -2 First term = -10 second term = a + d = -10 + (-2) = -10-2 = -12 Third term = a+ 2d = .......

OpenStudy (anonymous):

First 10 terms = -10,-12,-14,-16,-18,-20,-22,-24,-26,-28